OpenStudy (anonymous):

The period of vertical oscillation of a mass-spring system is T when the spring carries a mass of 1.00 kg. What mass should be added to the 1.00kg if the period was to be increased to 1.50T

OpenStudy (irishboy123):

the governing equation is \(m \ddot x + k x = 0\) so work from \(\omega = \sqrt{\dfrac{k}{m}}\)